OpenStudy (anonymous):

__/-- 20 (simplify the Expression) ( Radical)

OpenStudy (asnaseer):

your question does not seem to make sense - have you missed out some of the information?

OpenStudy (anonymous):

__/-- <<-- Radical Sign, __/--20 simplify it? thats what my book says

OpenStudy (asnaseer):

oh I see, so you want to simplify:\[\sqrt{20}\]

OpenStudy (anonymous):

how'd you do that?

OpenStudy (asnaseer):

you can type latex commands in the reply. I have added some tips on how to use latex in the latex practising group - you can join that and have a look in the closed questions section for tips

OpenStudy (anonymous):

Sure thanks! so the answer to my question is?

OpenStudy (asnaseer):

ok - first think of how you can write 20 as the the product of two numbers - one of which is a perfect square

OpenStudy (anonymous):

10 x 2?

OpenStudy (asnaseer):

neither 10 nor 2 is a perfect square - try again

OpenStudy (asnaseer):

numbers that are perfect squares can be written as \(x^2\) - e.g. 9 is a perfect square because it can be written as \(3^2\)

OpenStudy (anonymous):

20 x1? 5x 4?

OpenStudy (asnaseer):

5 x 4 is the correct answer because 4 is a perfect square - i.e. \(4=2^2\)

OpenStudy (anonymous):

so the answer is \[5\sqrt{4}\]

OpenStudy (asnaseer):

so now, we rewrite the expression as:\[\sqrt{20}=\sqrt{4*5}=\sqrt{4}*\sqrt{5}=2*\sqrt{5}=2\sqrt{5}\]

OpenStudy (asnaseer):

does that make sense?

OpenStudy (anonymous):

are you Simplifying the \[5\sqrt{4}\]

OpenStudy (asnaseer):

no - your answer is incorrect. I was simplifying \(\sqrt{20}\)

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

ty

OpenStudy (asnaseer):

yw

OpenStudy (anonymous):

how do i do Radical signs on the -Ask a question part-

OpenStudy (asnaseer):

join the Latex Practising Group - its all described there

OpenStudy (asnaseer):

look under the Closed Questions

OpenStudy (anonymous):

i tried to find that group but nothing shows up

OpenStudy (asnaseer):

^^You the link I just pasted

OpenStudy (asnaseer):

*Use

OpenStudy (anonymous):

were can i find it at? im on the latex page

OpenStudy (asnaseer):

look for how to enter inline latex

OpenStudy (asnaseer):

under the Closed Questions